OpenStudy (anonymous):

4b+6=-2 (PLEASE SHOW WORK)

OpenStudy (anonymous):

4b = -8 b = -8/4 b = -2

OpenStudy (anonymous):

4b +6 = -2 4b = -2 - 6 4b = -8 4b/4 = -8/4 b = -2